OpenStudy (anonymous):

please help! thank you :) one of the wavelengths of light emitted by hydrogen atoms is 656 nm, calculate the energy of a photon a. 4.57 x 10^14J b. 1.01 x 10^-27j c. 1.99 x 10^-25J d. 3.03 x 10^-19j Given the noble gas electron configuration below, identify the atom - [Kr] 5s^2 4d^2 a. Sn B. Sr c. Se d. Mo e. Zr what is the atom represented by the electron configuration 1s^2 2s^2 2p^6 3s^2 3p^6 4s^2 3d^4 a. Cr b.Mo c.O D.S E.Se

OpenStudy (anonymous):

also, if you don't mind i have a few more, which of the following orbitals does not exist? A. 7s B. 2d C.4f D.3p What is the noble gas electron configuration for Beryillium? A. [he] 2s^2 B.[Li] 2s^2 C. [Li] 2s^1 D. [He] 2s^1 Which color of light has a higher frequency? Red 780nm yellow 580nm blue 450nm green 530 Which of the following electron configurations represents a stable element? A. Na - 1s^2 2s^2 2p^6 3s^1 B. Cl - 1s^2 2s^2 2p^6 3s^2 3p^5 C. Ne - 1s^2 2s^2 2p^6 D. Li - 1s^2 2s^1 Determine the Electron configuration that corresponds to an atom in an excited state. A. 1s^2 2s^2 2p^6 3s^1 B. 1s^2 2s^2 2p^4 C.1s^2 2s^2 2p^5 3s^2 D.1s^2 2s^2 2p^6 which of the type of high energy electromagnetic radiation are likely to damage living organisms? Radio Infrared Visible Gamma please help, just a review, i really need to pass :) thank you ! :)

OpenStudy (anonymous):

E=hv Energy of a photon in joules= Planck's constant*freequency v=c/lambda frequency= speed of light/wavelength E (photon)= 6.63*10^-34 J-s *v v= 3.00*10^17 nm/ 656 nm Solve for v in the second equation and plug it into the second equation to determine energy.

OpenStudy (anonymous):

As for the orbital questions, this is a concept you'll have to become familiar with (using a periodic table). E. Zr A. Cr B. 2d A. [He] 2s^2

OpenStudy (anonymous):

I think you should be able to solve the rest yourself. Try watching a short video on youtube on orbitals in order to understand it well.

OpenStudy (anonymous):

Also, since it's not related to orbitals, I'll just tell you that it's gamma radiation.

OpenStudy (anonymous):

i honestly do not understand any of this, please just help me out, chemistry makes my head hurt

OpenStudy (anonymous):

i got D for the first question, i have no idea how to answer the rest

OpenStudy (anonymous):

is that right? anyone? please?